# Effective anisotropic interaction potentials for pairs of ultracold molecules shielded by a static electric field

### v1

@@ Abstract
Quantum gases of ultracold polar molecules have novel properties because of the strong dipolar forces between molecules. Current experiments shield the molecules from destructive collisions by engineering long-range repulsive interactions using microwave or static electric fields. These shielding methods produce interaction potentials with large repulsive cores that are not well described with contact potentials. In this paper we explore the anisotropic interaction potentials that arise for pairs of polar molecules shielded with static electric fields. We derive computationally inexpensive approximations for the potentials that are suitable for use in calculations of many-body properties. The interaction potentials for molecules shielded with static fields are qualitatively different from those that arise from single-field microwave shielding and will produce quite different many-body physics.
@@ I Introduction
To simulate many-body properties of systems made up of shielded ultracold molecules, we need simple representations of the effective potentials involved. For molecules shielded by a single microwave field [18], suitable potentials have been provided by Deng et al. [24] and used for many-body simulations [27]. In the microwave case the potentials are antidipolar at long range, meaning that they are repulsive for end-to-end approach of two molecules and attractive for side-to-side approach. At shorter range (inside a few hundred bohr), there is a repulsive potential that is itself strongly anisotropic.
For molecules shielded by static fields, the long-range potentials are qualitatively different from those for single-field microwave shielding. They are similar to those for dipoles fixed in space, and are attractive for end-to-end approach and repulsive for side-to-side approach. This is likely to lead to very different behaviour for many-body systems. Nevertheless, the effective potentials still have large repulsive cores [16, 17] and these are highly anisotropic. In this paper we investigate the anisotropic interaction potentials for molecules shielded by static fields and derive computationally inexpensive approximate forms for them that are suitable for quantum many-body calculations.
@@ II.1 Hamiltonian and its matrix elements
where $\mu_{\textrm{red}}$ is the reduced mass, ^$\hat{\boldsymbol{\nabla}}^{2}$ is the Laplacian, $\hat{h}_{k}$ is the Hamiltonian of molecule $k$ from Eq. 1, and $\hat{H}_{\textrm{dd}}$ represents the dipole-dipole interaction between the pair of molecules.
@@
where $\hat{\boldsymbol{r}}_{k}$ is a unit vector along the axis of molecule $k$. Here, the functions $C_{k,p}(\theta,\phi)$ are the Racah-normalised spherical harmonics and $\theta$ and $\phi$ indicate the orientation of $\boldsymbol{\hat{R}}$ with respect to the laboratory-frame axis $Z$. The tensor component $\left[\hat{\boldsymbol{r}}_{1}\otimes\hat{\boldsymbol{r}}_{2}\right]_{2,p}$ is
@@
×$\times C_{1,p_{1}}(\beta_{1},\alpha_{1})C_{1,p_{2}}(\beta_{2},\alpha_{2}).$ (5)
@@
×$\times\langle n_{2},m_{2}|C_{1,p_{2}}|n^{\prime}_{2},m^{\prime}_{2}\rangle C_{2,-p}(\theta,\phi),$ (6)
@@
×$\times\sqrt{(2n_{1}+1)(2n^{\prime}_{1}+1)}\begin{pmatrix}n_{1}&1&n^{\prime}_{1}\\ -m_{1}&p_{1}&m^{\prime}_{1}\end{pmatrix}\begin{pmatrix}n_{1}&1&n^{\prime}_{1}\\ 0&0&0\end{pmatrix}$
×$\times\sqrt{(2n_{2}+1)(2n^{\prime}_{2}+1)}\begin{pmatrix}n_{2}&1&n^{\prime}_{2}\\ -m_{2}&p_{2}&m^{\prime}_{2}\end{pmatrix}\begin{pmatrix}n_{2}&1&n^{\prime}_{2}\\ 0&0&0\end{pmatrix}.$ (7)
@@
×$\times\begin{pmatrix}n&1&n^{\prime}\\ -m&m-m^{\prime}&m^{\prime}\end{pmatrix}\begin{pmatrix}n&1&n^{\prime}\\ 0&0&0\end{pmatrix}.$ (9)
@@
×$\times\left(|\tilde{n}_{1},m_{1};\tilde{n}_{2},m_{2}\rangle\pm|\tilde{n}_{2},m_{2};\tilde{n}_{1},m_{1}\rangle\right).$ (11)
@@ II.2 Shielding with a static electric field
For collisions shielded by a static electric field, it is convenient to represent the field by a dimensionless quantity $\tilde{F}=F\mu/b$ [14], where $b$ is the rotational constant of the molecule. Similarly, the molecular energies are expressed in units of $b$. The pair state $(\tilde{n}_{1},m_{1})$+$(\tilde{n}_{2},m_{2})=(1,0)$+(1,0) crosses (0,0)+(2,0) at a critical electric field $\tilde{F}_{\textrm{X}}=3.244$, as shown in Fig. 1. At fields $\tilde{F}>\tilde{F}_{\textrm{X}}$, (1,0)+(1,0) lies above (0,0)+(2,0), and thus experiences repulsion due to interaction with it via $\hat{H}_{\textrm{dd}}$. This gives rise to shielding for the initial pair state (1,0)+(1,0).
@@
Figure 3: Contour plots of adiabats obtained from (a) full matrix diagonalisation, and (b) diagonalising the matrix of Eq. 12 for CaF at $F=23$ kV/cm. Contours are labeled in $\mu$K and shown only up to 50 $\mu$K.
@@ III.1 Perturbative model
The second term represents the repulsion due to the lower-lying level (0,0)+(2,0), while $V_{\textrm{rest}}(R,\theta)$ takes account of the interaction due to other pairs. In this model, both of these are treated perturbatively.
@@
This is smaller than the first two terms of Eq. III.1 at most $\theta$, except near $\theta=\theta_{0}$, where both $H_{11}$ and $H_{12}$ vanish. In the perturbative model, for simplicity, we therefore evaluate $V_{\textrm{rest}}$ only at $\theta=\theta_{0}$ and use the result at all $\theta$. At $\theta=\theta_{0}$, pair states with $m_{1}^{\prime}+m_{2}^{\prime}=0$ make no contribution and $V_{\textrm{rest}}(R)$ is dominated by a repulsive interaction due to (0,0)+(2,$\pm$1), which is the pair state that lies closest to (and below) (0,0)+(2,0) at these fields. All the state separations are proportional to the rotational constant $b$ except for a weak dependence on $R$. With this approximation,
@@
The dimensionless function $\alpha(\tilde{F})$ is shown in Fig. 6. We approximate it as a quadratic function of $F$,
$\alpha(\tilde{F})$ $\approx\alpha^{(0)}+\alpha^{(1)}(\tilde{F}-\tilde{F}_{\textrm{X}})+\alpha^{(2)}(\tilde{F}-\tilde{F}_{\textrm{X}})^{2},$ (19)
@@
with $\Delta(\tilde{F})$, $d_{ij}(\tilde{F})$ and $\alpha(\tilde{F})$ from Eqs. 15, 16 and 19, respectively.
Figure 8: Contour plot comparing $V^{\textrm{pert}}_{\textrm{eff}}$ (black lines) with adiabats obtained from full matrix diagonalisation (grey lines) for CaF at $F=23$ kV/cm. Contours are labeled in $\mu$K and shown only up to 50 $\mu$K.
@@
The perturbative model reproduces the qualitative features of the adiabats well. It corrects the main deficiency of Fig. 3(b) without significant extra computational expense. However, it produces some quantitative errors that may be significant in precise work. Its main deficiency is that it approximates the effect of pair states with $m_{1}^{\prime}+m_{2}^{\prime}\neq 0$ as isotropic and neglects entirely the effects of pair states other than (0,0)+(2,0) with $m_{1}^{\prime}+m_{2}^{\prime}=0$. This approximation becomes more serious at larger $\tilde{F}$, where the effects of (0,0)+(2,0) are not so strongly dominant. As seen in Fig.  7, it can cause underestimates of up to 20% in well depths at higher fields. It also starts to break down at small distances when the off-diagonal matrix element $H_{12}$ in Eq. 12 is comparable to the energy separation $\Delta$ between pair states 1 and 2. However, this occurs only for fields very close to $F_{\textrm{X}}$ and is most significant for molecules with small values of $\tilde{b}$. It may be seen in the lowest-field panels in Fig. 7 (for $\tilde{F}=3.25$); the repulsion at angles different from $\theta_{0}$ is overestimated for RbCs at 2.584 kV/cm but is reasonably accurate for CaF at 21.6 kV/cm.
@@ III.2 Fully coupled model
$Q_{12}^{(2)}$ $-5.19\times 10^{-1}$
$R_{11}^{(0)}$ $-7.36\times 10^{-2}$
$R_{11}^{(1)}$ $-3.36\times 10^{-2}$
@@
$R_{22}^{(1)}$ $-1.90\times 10^{-2}$
$R_{12}^{(0)}$ $-4.44\times 10^{-2}$
$R_{12}^{(1)}$ $-3.08\times 10^{-3}$
@@
$H_{ij,\textrm{VV}}(R,\theta,\tilde{F})$ $=\frac{\mu^{4}}{(4\pi\epsilon_{0})^{2}R^{6}}\frac{1}{b}\bigg{[}P_{ij}(\tilde{F})(1-3\cos^{2}\theta)^{2}$
@@
with $(i,j)=(1,1),(2,2)$ and $(1,2)$. The first term in square brackets represents second-order matrix elements of $H_{\textrm{dd}}$ between (1,0)+(1,0) and pair states in class 2 with $|m_{1}^{\prime}+m_{2}^{\prime}|=0$. These are proportional to the square of the spherical harmonics $|C_{2,0}(\theta,\phi)|^{2}$. Similarly, the second and third terms represent second-order interactions with class 2 levels with $|m_{1}^{\prime}+m_{2}^{\prime}|=1$ and 2, respectively, and are proportional to $|C_{2,\pm 1}(\theta,\phi)|^{2}$ and $|C_{2,\pm 2}(\theta,\phi)|^{2}$. The coefficients $P(\tilde{F})$, $Q(\tilde{F})$ and $R(\tilde{F})$ are dimensionless functions of field, and are approximated as
$P_{ij}(\tilde{F})$ $\approx P^{(0)}_{ij}+P^{(1)}_{ij}(\tilde{F}-\tilde{F}_{\textrm{X}});$ (25a)
$Q_{ij}(\tilde{F})$ $\approx Q^{(0)}_{ij}+Q^{(1)}_{ij}(\tilde{F}-\tilde{F}_{\textrm{X}})+Q^{(2)}_{ij}(\tilde{F}-\tilde{F}_{\textrm{X}})^{2};$ (25b)
$R_{ij}(\tilde{F})$ $\approx R^{(0)}_{ij}+R^{(1)}_{ij}(\tilde{F}-\tilde{F}_{\textrm{X}})$ (25c)
@@
Figure 10: Contour plot comparing $V^{\textrm{fc}}_{\textrm{eff}}$ (black lines) with adiabats obtained from full matrix diagonalisation (grey lines) for CaF at $F=23$ kV/cm. Contours are labeled in $\mu$K and shown only up to 50 $\mu$K.
@@ IV Contrast with microwave shielding
Figure 11: Contour plot of effective potential for shielding of CaF with circularly polarised microwaves at Rabi frequency $\Omega=30$ MHz and detuning $\Delta=0$, using the functional form of ref. [24]. Contours are labeled in $\mu$K and shown only up to 50 $\mu$K.
@@
The effective potentials for static-field shielding may be contrasted with those for microwave shielding with a circularly polarised field [24]. The range and depth of the potentials depend strongly on microwave Rabi frequency and detuning, but examples are shown in Fig. 11 for conditions that are reasonable for CaF. The potentials are attractive at long range for angles $\theta_{0}<\theta<180^{\circ}-\theta_{0}$, but repulsive at all distances for $\theta<\theta_{0}$. This is opposite to the situation for static-field shielding. Both static-field and microwave shielding provide large repulsive cores that are very different from the short-range (contact-like) interactions that are common in atomic systems. However, the dipole-dipole forces resulting from static-field shielding favour geometries around $\theta=0$, while those from microwave shielding with circular polarisation favour geometries around $\theta=90^{\circ}$.
@@
The radically different form of the effective potentials for static-field and microwave shielding will result in markedly different many-body physics. This will be true both for Bose-Einstein condensates of polar molecules and for polar molecules in optical lattices or tweezer arrays. For a sufficiently large number of particles and strong-enough Rabi coupling, microwave shielding is expected to result in a dense self-bound disc-like droplet [27, 28]. For large couplings, this may eventually undergo a transition into a crystalline monolayer of molecules [29]. In contrast, condensates of molecules shielded with static fields should (under proper conditions) form elongated filament-like droplets aligned along the direction of the field. These droplets, though resembling those observed in magnetic atoms [34], should not demand the stabilizing effect of quantum fluctuations. The properties of these self-bound droplets, including their stability, condensate fraction, and two-body correlations, are likely to be significantly different from those formed in the presence of microwave shielding. Condensates shielded with static fields may also present equilibrium multi-droplet solutions, which are not expected with microwave shielding [27, 29]. This will open intriguing questions about supersolidity [35] and pattern formation [36, 37, 38].
@@ V Conclusions
There is an important contrast between different types of shielding. Microwave shielding with circularly polarized fields gives effective potentials that are attractive for side-by-side geometries and favor molecules arranged in two-dimensional sheets. Static-field shielding, by contrast, gives effective potentials that are attractive for end-to-end geometries and favor molecules in linear arrangements. Microwave and static-field shielding are thus complementary, and both are needed to explore the full range of many-body behavior that can exist.
@@
The data that support the findings of this article are openly available from Durham University [at URL to be supplied later]. A subroutine that evaluates the effective potentials efficiently is available from the authors on request.

### v2

@@ Abstract
Quantum gases of ultracold polar molecules have novel properties because of the strong dipolar forces between molecules. Current experiments shield the molecules from destructive collisions by engineering long-range repulsive interactions using microwave or static electric fields. These shielding methods produce interaction potentials with large repulsive cores that are not well described with contact potentials. In this paper we explore the anisotropic interaction potentials that arise for pairs of polar molecules shielded with static electric fields. We derive computationally inexpensive approximations for the potentials that are suitable for use in calculations of many-body properties. The interaction potentials for molecules shielded with static fields are substantially different from those that arise from microwave shielding and will produce quite different many-body physics.
@@ I Introduction
To simulate many-body properties of systems made up of shielded ultracold molecules, we need simple representations of the effective potentials involved. For molecules shielded by a circularly polarised microwave field [18], suitable potentials have been provided by Deng et al. [24] and used for many-body simulations [27, 28]. In this case the potentials are antidipolar at long range, meaning that they are repulsive for end-to-end approach of two molecules and attractive for side-to-side approach. At shorter range (inside a few hundred bohr), there is a repulsive potential that is itself strongly anisotropic.
For molecules shielded by static fields, the long-range potentials are qualitatively different from those for circularly polarised microwaves. They are similar to those for dipoles fixed in space, and are attractive for end-to-end approach and repulsive for side-to-side approach. Nevertheless, the effective potentials still have large repulsive cores [16, 17] and these are highly anisotropic. Similar topologies can be achieved for microwave shielding by supplementing the circularly polarised field with a linearly polarised one [30, 31]. Such topologies are likely to lead to very different behaviour for many-body systems. In this paper we investigate the anisotropic interaction potentials for molecules shielded by static fields and derive computationally inexpensive approximate forms for them that are suitable for quantum many-body calculations.
@@ II.1 Hamiltonian and its matrix elements
where $\mu_{\textrm{red}}$ is the reduced mass, $\hat{\boldsymbol{\nabla}}^{2}$ is the Laplacian, $\hat{h}_{k}$ is the Hamiltonian of molecule $k$ from Eq. 1, and $\hat{H}_{\textrm{dd}}$ represents the dipole-dipole interaction between the pair of molecules.
@@
where $\hat{\boldsymbol{r}}_{k}$ is a unit vector along the axis of molecule $k$. Here, the functions $C_{k,p}(\theta,\phi)$ are the Racah-normalised spherical harmonics, and $\theta$ and $\phi$ indicate the orientation of $\boldsymbol{\hat{R}}$ with respect to the laboratory-frame axis $Z$. The tensor component $\left[\hat{\boldsymbol{r}}_{1}\otimes\hat{\boldsymbol{r}}_{2}\right]_{2,p}$ is
@@
$\times C_{1,p_{1}}(\beta_{1},\alpha_{1})C_{1,p_{2}}(\beta_{2},\alpha_{2}).$ (5)
@@
$\times\langle n_{2},m_{2}|C_{1,p_{2}}|n^{\prime}_{2},m^{\prime}_{2}\rangle C_{2,-p}(\theta,\phi),$ (6)
@@
$\times\sqrt{(2n_{1}+1)(2n^{\prime}_{1}+1)}\begin{pmatrix}n_{1}&1&n^{\prime}_{1}\\ -m_{1}&p_{1}&m^{\prime}_{1}\end{pmatrix}\begin{pmatrix}n_{1}&1&n^{\prime}_{1}\\ 0&0&0\end{pmatrix}$
$\times\sqrt{(2n_{2}+1)(2n^{\prime}_{2}+1)}\begin{pmatrix}n_{2}&1&n^{\prime}_{2}\\ -m_{2}&p_{2}&m^{\prime}_{2}\end{pmatrix}\begin{pmatrix}n_{2}&1&n^{\prime}_{2}\\ 0&0&0\end{pmatrix}$. (7)
@@
$\times\begin{pmatrix}n&1&n^{\prime}\\ -m&m-m^{\prime}&m^{\prime}\end{pmatrix}\begin{pmatrix}n&1&n^{\prime}\\ 0&0&0\end{pmatrix}$. (9)
@@
$\times\left(|\tilde{n}_{1},m_{1};\tilde{n}_{2},m_{2}\rangle\pm|\tilde{n}_{2},m_{2};\tilde{n}_{1},m_{1}\rangle\right).$ (11)
@@ II.2 Shielding with a static electric field
For collisions shielded by a static electric field, it is convenient to represent the field by a dimensionless quantity $\tilde{F}=F\mu/b$ [14], where $b$ is the rotational constant of the molecule. Similarly, the molecular energies are expressed in units of $b$. The pair state $(\tilde{n}_{1},m_{1})$$(\tilde{n}_{2},m_{2})=(1,0)$+(1,0) crosses (0,0)+(2,0) at a critical electric field $\tilde{F}_{\textrm{X}}=3.244$, as shown in Fig. 1. At fields $\tilde{F}>\tilde{F}_{\textrm{X}}$, (1,0)+(1,0) lies above (0,0)+(2,0), and thus experiences repulsion due to interaction with it via $\hat{H}_{\textrm{dd}}$. This gives rise to shielding for the initial pair state (1,0)+(1,0).
@@
[…]
@@ III.1 Perturbative model
The second term represents the repulsion due to the lower-lying pair state (0,0)+(2,0), while $V_{\textrm{rest}}(R,\theta)$ takes account of the interaction due to other pairs. In this model, both of these are treated perturbatively.
@@
[…]
@@
[…]
@@
[…]
@@
The perturbative model reproduces the qualitative features of the adiabats well. It corrects the main deficiency of Fig. 3(b) without significant extra computational expense. However, it produces some quantitative errors that may be significant in precise work. Its main deficiency is that it approximates the effect of pair states with $m_{1}^{\prime}+m_{2}^{\prime}\neq 0$ as isotropic, as in Eq. 18. Furthermore, it neglects the effects of pair states other than (0,0)+(2,0) with $m_{1}^{\prime}+m_{2}^{\prime}=0$ entirely, because Eq. 18 is evaluated at $\theta_{0}$, where $P_{2}(\cos\theta)=0$. The approximation becomes more serious at larger $\tilde{F}$, where the effects of (0,0)+(2,0) are not so strongly dominant. As seen in Fig. 7, it can cause underestimates of up to 20% in well depths at higher fields. It also starts to break down at small distances when the off-diagonal matrix element $H_{12}$ in Eq. 12 is comparable to the energy separation $\Delta$ between pair states 1 and 2. However, this occurs only for fields very close to $F_{\textrm{X}}$ and is most significant for molecules with small values of $\tilde{b}$. It may be seen in the lowest-field panels in Fig. 7 (for $\tilde{F}=3.25$); the repulsion at angles different from $\theta_{0}$ is overestimated for RbCs at 2.584 kV/cm but is reasonably accurate for CaF at 21.6 kV/cm.
@@ III.2 Fully coupled model
$Q_{12}^{(2)}$ $5.19\times 10^{-1}$
$R_{11}^{(0)}$ $7.36\times 10^{-2}$
$R_{11}^{(1)}$ $3.36\times 10^{-2}$
@@
$R_{22}^{(1)}$ $1.90\times 10^{-2}$
$R_{12}^{(0)}$ $4.44\times 10^{-2}$
$R_{12}^{(1)}$ $3.08\times 10^{-3}$
@@
[…]
@@
with $(i,j)=(1,1),(2,2)$ and $(1,2)$. The first term in square brackets represents second-order matrix elements of $H_{\textrm{dd}}$ between (1,0)+(1,0) and pair states in class 2 with $|m_{1}^{\prime}+m_{2}^{\prime}|=0$. These are proportional to the square of the spherical harmonics $|C_{2,0}(\theta,\phi)|^{2}$. Similarly, the second and third terms represent second-order interactions with pair states in class 2 with $|m_{1}^{\prime}+m_{2}^{\prime}|=1$ and 2, respectively, and are proportional to $|C_{2,\pm 1}(\theta,\phi)|^{2}$ and $|C_{2,\pm 2}(\theta,\phi)|^{2}$. The coefficients P(F~)P($\tilde${F}), Q(F~)Q($\tilde${F}) and R(F~)R($\tilde${F}) are dimensionless functions of field, and are approximated as
[…]
@@
[…]
@@ IV Contrast with microwave shielding
[…]
@@
The effective potentials for static-field shielding may be contrasted with those for microwave shielding with a circularly polarised field [24]. The range and depth of the latter depend strongly on the microwave Rabi frequency and detuning. An example is shown in Fig. 11 for conditions that are reasonable for CaF. The potential is attractive at long range for angles $\theta_{0}<\theta<180^{\circ}-\theta_{0}$, but repulsive at all distances for $\theta<\theta_{0}$. This is opposite to the situation for static-field shielding. Both static-field and microwave shielding provide large repulsive cores that are very different from the short-range (contact-like) interactions that are common in atomic systems. However, the dipole-dipole forces resulting from static-field shielding favour geometries around $\theta=0$, while those from microwave shielding with circular polarisation favour geometries around $\theta=90^{\circ}$.
Interaction potentials with minima around $\theta=0$ can also be produced by double microwave shielding, combining circularly and linearly polarised microwave fields [30, 31]. In this case, however, shielding is maintained only when circular polarisation dominates. Strong losses occur with linear polarisation alone [18], and this limits the strength of dipolar (as opposed to antidipolar) interactions that can be achieved with microwave shielding. Fig. 12 shows an example of the dipolar interaction potential achievable with double microwave shielding for CaF. The potential is considerably weaker and has a different shape to that obtained with static-field shielding. Interaction potentials with minima around a single direction in space can also be produced by shielding with elliptical polarisation [36]. However, the resulting interaction potentials are noncylindrical and depend on a large parameter space of microwave parameters, making detailed comparisons complicated. Moreover, elliptical polarisation causes enhanced losses [20].
A further difference between microwave and static-field shielding is that microwave shielding is near-universal, in the sense that the effective potentials are similar for all molecules when expressed in terms of the dipole length $R_{3}$ and energy $E_{3}$ [37]. With static fields, by contrast, the effective potentials depend strongly on $\tilde{b}$, as described in Sec. II.2: they are deepest and strongest (in reduced units) for molecules with large $\tilde{b}$ [17].
@@
Effective potentials with attractive interactions around $\theta=0$ and $90^{\circ}$ will result in markedly different many-body physics. This will be true both for Bose-Einstein condensates of polar molecules and for polar molecules in optical lattices or tweezer arrays. For a sufficiently large number of molecules and strong-enough Rabi coupling, attraction around $90^{\circ}$ is expected to result in a dense self-bound disc-like droplet [27, 28]. For large couplings, this may eventually undergo a transition into a crystalline monolayer of molecules [29]. In contrast, molecular condensates with attraction around $\theta=0$ should (under proper conditions) form elongated filament-like droplets aligned along the direction of the field. These droplets, though resembling those observed in magnetic atoms [38], should not demand the stabilizing effect of quantum fluctuations. The properties of these self-bound droplets, including their stability, condensate fraction, and two-body correlations, are likely to be significantly different from those formed in the presence of microwave shielding. They may also present equilibrium multi-droplet solutions, which are not expected when the attraction is localised around $90^{\circ}$ [27, 29]. This will open intriguing questions about supersolidity [39] and pattern formation [40, 41, 42].
@@ V Conclusions
There are important contrasts between different types of shielding. Microwave shielding with circularly polarized fields gives effective potentials that are attractive for side-by-side geometries and favor molecules arranged in two-dimensional sheets. Static-field shielding, by contrast, gives effective potentials that are attractive for end-to-end geometries and favor molecules in linear arrangements. Microwave shielding with combined circular and linear polarisations can also give effective potentials that favour linear arrangements, but with important differences in strength and shape. Microwave and static-field shielding are complementary, and both are needed to explore the full range of many-body behavior that can exist.
@@
The data that support the findings of this article are openly available at https://doi.org/10.5281/zenodo.16966525. A subroutine that evaluates the effective potentials efficiently is available from the authors on request.